# Spacetime Spins: Statistical mechanics for error correction with stabilizer circuits

### v1

@@ III.2 Independent X-Z noise
We use the swapped subscripts on $K_{X}$ and $K_{Z}$ to increase interpretability: $K_{\alpha}$ interaction strengths account for the probability of $\alpha$-errors. Hamiltonian to separate into $X$ and $Z$ contributions. The Hamiltonian’s interaction signs are sampled from independent Bernoulli distributions, where
@@
for every qubit $i$ and $\alpha=X,Z$. Ignoring the constant contribution to the Hamiltonian from the $\alpha=I$ terms, Eq. (14) is then written as
@@ III.3 Spin diagrams
One of our central results is to provide a framework to systematically construct these Hamiltonians using spin diagrams: a graphical language where each circuit element—idle wires, CNOTs, measurements, etc.—contributes standard “building blocks” of spins and interactions. These diagrams both enable constructing models for arbitrary Clifford circuits, and reveal connections to established lattice models. Spin diagrams work for general noise channels; here we demonstrate them for independent circuit-level $X$-$Z$ noise, where they are particularly transparent. For general noise channels, each spin has more interactions, which are harder to visualize.
We construct each diagram by drawing a square for each spin $\sigma_{k}$, and connect red (blue) lines for each field term, bond, or many-body interaction in $H_{X,E}$ ($H_{Z,E}$). The number of lines attached to a spin are the number of interactions it participates in through the Hamiltonian—we refer to this as the degree of the spin. Figure 1 illustrates some of these building blocks. An idling qubit, for example, contributes two generators and therefore two spins, one for $H_{X}$ (red) and one for $H_{Z}$ (blue). A CNOT gate couples together spins on adjacent qubits. Composing these building blocks together, we can represent a circuit as a spin diagram, such as in Fig. 2.
@@
Many of these spins can be eliminated, producing a simpler—but equivalent—Hamiltonian: Spins with a single interaction, arising around single-qubit resets and measurements, correspond to gauge operators that identify single-qubit errors with no effect on the syndrome or logical operators. Spins with exactly two interactions, such as those found on idling qubits, encode gauge-equivalent pairs of single-qubit errors at different spacetime locations that can be grouped together. Both cases can be integrated out, simplifying the model while preserving the partition function up to an $E$-independent multiplicative factor; relevant physical quantities like probability ratios $\mathbb{P}(\overline{E})/\mathbb{P}(\overline{E}^{\prime})$, free-energy differences $\Delta F=-\ln\mathcal{Z}_{E}+\ln\mathcal{Z}_{E^{\prime}}$, and thus ML decoding decisions remain unchanged. To preserve the partition function, each interaction must now be assigned a weight equal to the number of single-qubit errors grouped together as gauge-equivalent. This weight determines the interaction’s effective coupling strength $K$ and the probability distribution of its sign $\eta$, see Eqs. (25) and (26) in Sec. III.4, with additional details in Appendix A.2.1. We illustrate this simplification on a spin diagram corresponding to the $ZZ$ measurement circuit in Fig. 2.
@@ III.4 Gauge-equivalent components
Steps 1 and 2 respectively carry out the process of removing order-$1$ spins (corresponding to single-qubit generators; we can remove these spins as this preserves the partition function up to a constant multiplicative factor) and grouping together interactions for order-$2$ spins. In this way, each connected component includes interactions from the unintegrated Hamiltonian that are combined together into one weighted interaction in the integrated Hamiltonian. Some connected components may still consist of only one vertex: these are interactions from the Hamiltonian that persist after integration.
Prior to integration, each interaction represents the probability of an error occurring at a spacetime location and with a Pauli flavor $\alpha=X,Z$. To preserve the partition function, the new interactions must take into account the probabilities of errors occurring at any of the spacetime locations in each of the connected components. Only the parity of the number of errors is important: by construction, an even number of errors on a connected component is gauge-equivalent to the identity. If a connected component includes $x$ vertices labeled $\alpha=X$, and $z$ for $\alpha=Z$, the probability that an odd total number of errors occurs across that component is
@@
$\Xi_{c}=\left\{k:\left\llbracket{\prod_{(\alpha,i,\tau)\in c}[\tilde{\alpha}]_{i,\tau}},{g_{k}}\right\rrbracket=-1\right\},$ (28)
@@
For single-vertex connected components, $p_{\alpha}^{(1)}=p_{\alpha}$ and $K_{\alpha}^{(1)}=K_{\alpha}$. As the size of a connected component increases, $p_{X}^{(|c|_{x})}\rightarrow\frac{1}{2}$ and $K_{X}^{(|c|_{X})}\rightarrow 0$: physically, a net odd-parity error is more likely to occur on larger components, and consequently the Hamiltonian places a smaller energy penalty on that interaction being violated. If the size of a component scales extensively with the number of spacetime qubits, in the thermodynamic limit its interaction becomes irrelevant to the free energy.
@@ IV Repetition code
We now illustrate our framework through three simple examples, each using the 1D repetition code. First, we derive the spin models for a memory experiment and a stability experiment on the code, using a simple measurement-only brickwork circuit that elucidates the spacetime duality of these two diagnostic tools. Secondly, we compare different syndrome-extraction circuits, each mapping onto random-bond Ising models with distinct bond anisotropies. Thirdly, we examine transversal CNOT gates between repetition code patches, showing how logical operations introduce defects into these models. In these examples, we estimate maximum-likelihood decoding thresholds using Monte Carlo simulations and compare to minimum-weight perfect-matching (MWPM) decoders implemented with Stim [gidneyStim2021] and pymatching [higgottSparse2025]. Numerical methods are detailed in Appendix B.
@@ IV.1 Memory and stability experiments
Figure 4: Measurement circuits for repetition code experiments, where errors occur in-between each circuit operation. We show representatives of error cosets that lead to decoding failures. The detector cells triggered by these errors are highlighted in yellow; some untriggered detector cells are shown in pink. Both circuits map onto a random-bond Ising model on a rotated square lattice. The errors flip the sign of bonds, indicated by dashed lines, and form a chain on the dual lattice. Detector cells form plaquettes of the lattice, and triggered detector cells host Ising vortices at the endpoints of the chains. The spatial and temporal (open) boundary conditions are interchanged between the two experiments.
Syndrome-extraction circuits can be used for different experimental tasks; two common diagnostic tools are the memory and stability experiments (cf. Sec. III.1). Notably, the two experiments can be interpreted as spacetime duals of each other [gidneyStability2022, liPerturbative2025]. As we explain in this section, this is clearly seen using their spacetime spin diagrams. The structure of these diagrams gives insight to the differences (or lack thereof) between memory and stability experimental thresholds.
We focus on the case where two-qubit $ZZ$ measurements are the elementary building blocks. Taking into account that each qubit may undergo only one measurement at a time, the circuit takes on a brickwork spatiotemporal layout, shown in Fig. 4. (This already describes a circuit-level model in systems with native pairwise measurements [grans-samuelssonImproved2024, grans-samuelssonFaulttolerant2025].)
@@
It is well-established that the repetition code subjected to measurement errors under a phenomenological noise model maps onto the square lattice RBIM [dennisTopological2002]. Here, spin diagrams have allowed us to easily derive the associated spin model for our circuit: an RBIM, but on a rotated square lattice. Unlike previous works, we have no explicit measurement errors; the non-simultaneous layering of gates (i.e., brickwork) and the repeated syndrome measurements bring about the two-dimensional structure.
@@
Figure 4b shows a stability experiment for the repetition code, again using only measurements. The initialization and final measurements are now in the $X$ basis to prevent detector cells from forming on the temporal boundaries, enabling logical equivalence classes with undetected spanning timelike errors, while partial stabilizers [^8] These are logical $Z$ operators for the repetition code, hence they remove the logical degree of freedom; equivalently, they define boundary stabilizers for the spacetime code on which $X$-strings cannot terminate. are measured on the spatial boundaries to detect spanning spacelike errors and remove these unwanted equivalence classes from the experiment [gidneyStability2022]. The bulk of the Hamiltonian is again a rotated RBIM. On the spatial boundaries, however, we now have field terms due to the single-qubit $Z$ measurements truncating the interactions from each $M_{ZZ}$ spin. On the temporal boundaries, weight-$2$ spatial bonds are created after removing order-$1$ spins from each $R_{X}$ and $M_{X}$, leaving the adjacent $M_{ZZ}$ spins as order-$2$.
@@
Figure 5: Error rates for memory and stability experiments corresponding to the repetition code measurement circuits in Figs. 4a and 4b respectively, using MWPM decoders with circuits of size $N=d$, $T=2d+1$. Both exhibit the same threshold, the noise rate $p_{X}$ below which increasing $d$ decreases the experimental error rate: approximately $10.0\%$.
@@
From the spin diagrams, the spacetime duality between the memory experiment and stability experiments is apparent. The error thresholds of the repetition code in a memory experiment (determined by the free energy cost of a spacelike logical operator) and in a stability experiment (from the free energy cost of an undetected spanning timelike error) should be identical. To test this, we perform numerical simulations of the circuit using the MWPM decoder. Each run a syndrome is passed to the decoder, which estimates the underlying error. The memory experiment is a success if the decoder predicts an eigenvalue of $\bar{Z}=Z_{1}$ (based on anticommutation with its estimated error) that is consistent with the true eigenvalue at the end of the circuit. Similarly, the stability experiment is a success if the decoder predicts the correct value for the product of stabilizer measurement outcomes $Z_{1},\,Z_{1}Z_{2},\,Z_{2}Z_{3},\,\ldots,\,Z_{N-1}Z_{N},\,Z_{N}$. [^9] Qubit $1$ is on the left of Fig. 4b; $Z_{1}$ and $Z_{N}$ are the single-qubit $M_{Z}$ on the spatial boundaries. In the absence of errors, this product is always trivial (each $Z_{i}$ is included twice). In Fig. 4b, for example, the true outcome would be 1-1 due to the $X$ error before the final round of $M_{ZZ}$. Figure 5 shows the failure rates of these experiments at different bit-flip noise rates $p_{X}$. Vertical lines indicate the approximate location of the threshold computed via least-squares estimation in the vicinity of the crossing (cf. Appendix B.2). Both experiments, as predicted, exhibit an equal threshold of approximately $10.0\%$.
@@
The ordered-disordered phase transition of a square lattice RBIM along the Nishimori line is known to occur at $p=10.94(2)\%$ [honeckerUniversality2001, merzTwodimensional2002]. Accounting for boundaries—contributing field terms and enhanced finite-size effects—it is reasonable to expect a threshold to deviate from this value. Indeed, running Monte Carlo simulations (cf. Appendix B) to compute the free energy of the spin model for the memory experiment, Fig. 6 highlights a crossing point in the logical error rate at $p=10.77(8)\%$. MWPM decoders consider only the most-likely error rather than the error cosets and are therefore not optimal; the lower threshold of approximately $p_{X}=10\%$ and consistently higher logical error rates (at $p_{X}=10\%$, the MWPM exhibits a logical error rate of approximately $26\%$ compared to the SM Monte Carlo $24-25\%$, for example) reflect this. This discrepancy in MWPM threshold from the $10.5(2)\%$ zero-temperature critical value found in Ref. [kawashimaZeroTemperature1999] may also be accounted for by finite size effects.
@@ IV.2 Circuit compilations
The stabilizers of a QEC code can compiled into syndrome-extraction circuits in a variety of ways. The choice of implementation may be influenced by factors such as the native hardware-efficient gate set or a need to tolerate fabrication defects [debroyLUCI2024, wolanskiAutomated2025, mcewenRelaxing2023, eickbuschDemonstration2025]. In this section, we demonstrate how different syndrome-extraction circuits for the repetition code can be mapped onto spin diagrams with distinct bond anisotropies. We compare these models to offer insight into the differing performance of these circuits.
@@
mcewenRelaxing2023 introduced an alternate, dynamic syndrome-extraction circuit for the repetition code that exchanges the roles of data and ancilla qubits, producing a “wiggling” pattern in the detector cells. Figure 7b shows this circuit and the corresponding spin diagrams. As with the regular circuit, the simplified model (cf. Fig. 7d) is a random bond Ising model, this time on an un-sheared honeycomb lattice with anisotropic weights $w_{\braket{i,j}}\in\{2,2,3\}$. In the bulk of the model, the shear does not affect local interactions—the difference in memory-experiment threshold between these models should be solely due to the differing anisotropy.
Figure 8: MWPM error rates for memory experiments of the standard (solid line) and wiggling (dot-dashed line) repetition code circuits with N=2d1N=2d-1, T=4d1T=4d-1. The wiggling circuit has a lower threshold at $2.87(1)\%$ compared to $2.94(1)\%$ for the standard circuit.
To understand the difference in decoding performance of these two circuits, consider the free energy cost of inserting a spatial logical operator. One representative of this logical operator is the product of $X$ over all qubits in-between each measurement and reset, indicated by the circular errors and dashed bonds in Fig. 7. Starting from a ground state with $\eta_{X,\braket{i,j}}=1$ and $\sigma_{i}=1$ for both models, applying the logical error flips vertical weight-$3$ bonds. The energy cost is identical for both circuits. However, we need to account for the energies of gauge-equivalent errors when computing free-energy differences. Flipping a spin attached to the affected bonds (i.e., including a gauge operator to deform the logical error) changes energy by $\Delta\mathcal{E}_{\text{standard}}=2K_{X}^{(1)}$ in the standard circuit, and $\Delta\mathcal{E}_{\text{wiggling}}=4K_{X}^{(2)}-2K_{X}^{(3)}$ in the wiggling circuit. Taking the low-error limit of Eq. (25), one can show that
@@
We performed simulations of MWPM decoding for a memory experiment with both circuits. We simulated a circuit depth of T=4d1T=4d-1; below that, the code displayed above-threshold behavior with logical error rate increasing with distances. We plot the results of the memory experiment in Fig. 8: the standard circuit consistently achieves lower logical error rates, as we expected based on our spin models. The standard circuit has a MWPM threshold of approximately $2.94(1)\%$, while the wiggling circuit has a lower threshold of $2.87(1)\%$. This is also consistent with our expectations. Running Monte Carlo simulations of the spin models and estimating the ML decoding error rates, Fig. 9 shows a threshold of $2.98(1)\%$ for the standard circuit, slightly exceeding the MWPM threshold. The logical error rates themselves are also lower for the ML decoder, achieving a rate of $19\%$ at $p_{X}=2.7\%$ and $d=11$, compared to $22\%$ for the MWPM decoder, for example.
@@
Figure 9: Monte Carlo simulations of the logical error rate for a standard repetition code circuit with N=2d1N=2d-1, T=4d1T=4d-1. A threshold of $2.98(1)\%$ exceeds the MWPM threshold of $2.94(1)\%$ from Fig. 8.

### v2

@@ III.2 Independent X-Z noise
We use the swapped subscripts on $K_{X}$ and $K_{Z}$ to increase interpretability: $K_{\alpha}$ interaction strengths account for the probability of $\alpha$-errors for $\alpha=X,Z$. $Y$ errors manifest as $X$ and $Z$ errors on the same qubit and are not accounted for separately in the Hamiltonian. Interaction signs are sampled from independent Bernoulli distributions, where
@@
for every qubit $(i,\tau)$ and $\alpha=X,Z$. Ignoring the constant contribution to the Hamiltonian from the $\alpha=I$ terms, Eq. (14) is then written as
@@ III.3 Spin diagrams
One of our central results is to provide a framework to systematically construct these Hamiltonians using spin diagrams: a graphical language where each circuit element—idle wires, CNOTs, measurements, etc.—contributes standard “building blocks” of spins and interactions. These diagrams offer concise representations of all gauge-equivalent error configurations and their associated probabilities. Spin diagrams work for general noise channels; here we demonstrate them for independent circuit-level $X$-$Z$ noise, where they are particularly transparent. For general noise channels, each spin has more interactions, which are harder to visualize.
We construct a diagram by drawing a square for each spin $\sigma_{k}$, and connect red (blue) lines for each field term, bond, or many-body interaction in $H_{X,E}$ ($H_{Z,E}$). The number of lines attached to a spin are the number of interactions it participates in through the Hamiltonian—we refer to this as the degree of the spin. Figure 1 illustrates some of these building blocks. An idling qubit, for example, contributes two generators and therefore two spins, one for $H_{X}$ (red) and one for $H_{Z}$ (blue). A CNOT gate couples together spins on adjacent qubits. Composing these building blocks together, we can represent a circuit as a spin diagram, such as in Fig. 2.
@@
Many of these spins can be eliminated, producing a simpler—but equivalent—Hamiltonian: Spins with a single interaction, arising around single-qubit resets and measurements, correspond to gauge operators that identify single-qubit errors with no effect on the syndrome or logical operators (called “absolutely trivial fault configurations” in Ref. [15]). Spins with exactly two interactions, such as those found on idling qubits, encode gauge-equivalent pairs of single-qubit errors at different spacetime locations that can be grouped together. Both cases can be integrated out, simplifying the model while preserving the partition function up to an $E$-independent multiplicative factor; relevant physical quantities like probability ratios $\mathbb{P}(\overline{E})/\mathbb{P}(\overline{E}^{\prime})$, free-energy differences $\Delta F=-\ln\mathcal{Z}_{E}+\ln\mathcal{Z}_{E^{\prime}}$, and thus ML decoding decisions remain unchanged. To preserve the partition function, each interaction must now be assigned a weight equal to the number of single-qubit errors grouped together as gauge-equivalent. This weight determines the interaction’s effective coupling strength $K$ and the probability distribution of its sign $\eta$, see Eqs. (25) and (26) in Sec. III.4, with additional details in Appendix A.2.1. These rules are illustrated diagrammatically in Fig. 3 (an additional rule is discussed in Appendix A.3), which we then apply to simplify the example $ZZ$ measurement circuit from Fig. 2. This simplification reveals properties of the circuit; for example, $H_{Z}$ contains only one interaction as a $Z$ error on any of the spacetime locations is equivalent (the circuit only detects $X$ errors).
@@ III.4 Gauge-equivalent components
Steps 1 and 2 respectively carry out the process of removing degree-$1$ spins (corresponding to single-qubit generators; we can remove these spins as this preserves the partition function up to a constant multiplicative factor) and grouping together interactions for degree-$2$ spins (cf. Fig. 3). In this way, each connected component includes interactions from the unintegrated Hamiltonian that are combined together into one weighted interaction in the integrated Hamiltonian. Some connected components may still consist of only one vertex: these are interactions from the Hamiltonian that persist after integration.
Prior to integration, each interaction represented the probability of an error occurring at a spacetime location and with a Pauli flavor $\alpha=X,Z$. To preserve the partition function, the new interactions must take into account the probabilities of errors occurring at any of the spacetime locations in each of the connected components. Only the parity of the number of errors is important: by construction, an even number of errors on a connected component is gauge-equivalent to the identity. If a connected component includes $x$ vertices labeled $\alpha=X$, and $z$ for $\alpha=Z$, the probability that an odd total number of errors occurs across that component is
@@
$\Xi_{c}=\left\{k:\left\llbracket{\prod_{(\alpha,i,\tau)\in c}[\tilde{\alpha}]_{i,\tau}},{g_{k}}\right\rrbracket=-1\right\}$ (28)
@@
For single-vertex connected components, $p_{\alpha}^{(1)}=p_{\alpha}$ and $K_{\alpha}^{(1)}=K_{\alpha}$. As the size of a connected component increases, $p_{X}^{(w_{X,c})}\rightarrow\frac{1}{2}$ and $K_{X}^{(w_{X,c})}\rightarrow 0$: physically, a net odd-parity error is more likely to occur on larger components, and consequently the Hamiltonian places a smaller energy penalty on that interaction being violated. If the size of a component scales extensively with the number of spacetime qubits, in the thermodynamic limit its interaction becomes irrelevant to the free energy.
@@ IV Repetition code
We now illustrate our framework through three simple examples, each using the 1D repetition code. First, we derive the spin models for a memory experiment and a stability experiment on the code, using a simple measurement-only brickwork circuit that elucidates the spacetime duality of these two diagnostic tools. Secondly, we compare different syndrome-extraction circuits, each mapping onto random-bond Ising models with distinct bond anisotropies. Thirdly, we examine transversal CNOT gates between repetition code patches, showing how logical operations introduce defects into these models. In these examples, we estimate maximum-likelihood decoding thresholds using Monte Carlo simulations and compare to minimum-weight perfect-matching (MWPM) decoders implemented with Stim [54] and pymatching [63]. Numerical methods are detailed in Appendix B.
@@ IV.1 Memory and stability experiments
Figure 5: Brickwork measurement circuits for repetition code experiments, where errors occur in-between each circuit operation. We show representatives of error cosets that lead to decoding failures. The detector cells triggered by these errors are highlighted in yellow; some untriggered detector cells are shown in pink. Both circuits map onto a random-bond Ising model on a rotated square lattice. The errors flip the sign of bonds, indicated by dashed lines, and form a chain on the dual lattice. Detector cells form plaquettes of the lattice, and triggered detector cells host Ising vortices at the endpoints of the chains. The spatial and temporal (open) boundary conditions are interchanged between the two experiments.
Syndrome-extraction circuits can be used for different experimental tasks; two common diagnostic tools are the memory and stability experiments (cf. Sec. III.1). Notably, the two experiments can be interpreted as spacetime duals of each other [55, 77]. As we explain in this section, this is clearly seen using their spacetime spin diagrams. The structure of these diagrams gives insight to the differences (or lack thereof) between memory and stability experimental thresholds.
We use two-qubit $ZZ$ measurements as the elementary building blocks. Taking into account that each qubit may undergo only one measurement at a time, the circuit takes on a brickwork spatiotemporal layout, shown in Fig. 5. This already describes a circuit-level model in systems with native pairwise measurements [59, 58], and otherwise presents a phenomenological representation of a measurement circuit.
@@
It is well-established that the repetition code subjected to measurement errors under a phenomenological noise model maps onto the square lattice RBIM [37]. Here, spin diagrams have allowed us to easily derive the associated spin model for our circuit: an RBIM, but on a rotated square lattice. Unlike previous works, we have no explicit measurement errors; the non-simultaneous layering of gates (i.e., brickwork) and the repeated syndrome measurements bring about the two-dimensional structure.
@@
Figure 5b shows a stability experiment for the repetition code, again using only measurements. The initialization and final measurements are now in the $X$ basis to prevent detector cells from forming on the temporal boundaries, enabling logical equivalence classes with undetected spanning timelike errors, while partial stabilizers [^8] These are logical $Z$ operators for the repetition code, hence they remove the logical degree of freedom; equivalently, they define boundary stabilizers for the spacetime code on which $X$-strings cannot terminate. are measured on the spatial boundaries to detect spanning spacelike errors and remove these unwanted equivalence classes from the experiment [55]. The bulk of the Hamiltonian is again a rotated RBIM. On the spatial boundaries, however, we now have field terms due to the single-qubit $Z$ measurements truncating the interactions from each $M_{ZZ}$ spin. On the temporal boundaries, weight-$2$ spatial bonds are created after removing degree-$1$ spins from each $R_{X}$ and $M_{X}$, leaving the adjacent $M_{ZZ}$ spins as degree-$2$.
@@
Figure 6: Error rates for memory and stability experiments corresponding to the repetition code measurement circuits in Figs. 5a and 5b respectively, using MWPM decoders with circuits of size $N=d$, $T=2d+1$. Both exhibit the same threshold, the noise rate $p_{X}$ below which increasing $d$ decreases the experimental error rate: approximately $10.0\%$.
@@
From the spin diagrams, the spacetime duality between the memory experiment and stability experiments is apparent. The error thresholds of the repetition code in a memory experiment (determined by the free energy cost of a spacelike logical operator) and in a stability experiment (from the free energy cost of an undetected spanning timelike error) should be identical. To test this, we perform numerical simulations of the circuit using the MWPM decoder. Each run, a syndrome is passed to the decoder, which estimates the underlying error. The memory experiment is a success if the decoder predicts an eigenvalue of $\bar{Z}=Z_{1}$ (based on anticommutation with its estimated error) that is consistent with the true eigenvalue at the end of the circuit. Similarly, the stability experiment is a success if the decoder predicts the correct value for the product of stabilizer measurement outcomes $Z_{1},\,Z_{1}Z_{2},\,Z_{2}Z_{3},\,\ldots,\,Z_{N-1}Z_{N},\,Z_{N}$. [^9] Qubit $1$ is on the left of Fig. 5b; $Z_{1}$ and $Z_{N}$ are the single-qubit $M_{Z}$ on the spatial boundaries. In the absence of errors, this product is always trivial (each $Z_{i}$ is included twice). In Fig. 5b, for example, the true outcome would be 1-1 due to the $X$ error before the final round of $M_{ZZ}$. Figure 6 shows the failure rates of these experiments at different bit-flip noise rates $p_{X}$. Vertical lines indicate the approximate location of the threshold computed via least-squares estimation in the vicinity of the crossing (cf. Appendix B.2). Both experiments, as predicted, exhibit an equal threshold of approximately $10.0\%$.
@@
The ordered-disordered phase transition of a square lattice RBIM along the Nishimori line is known to occur at $p=10.94(2)\%$ [64, 86]. Accounting for boundaries—contributing field terms and enhanced finite-size effects—it is reasonable to expect a threshold to deviate from this value. Indeed, running Monte Carlo simulations (cf. Appendix B) to compute the free energy of the spin model for the memory experiment, Fig. 7 highlights a crossing point in the logical error rate at $p=10.77(8)\%$. MWPM decoders consider only the most-likely error rather than the error cosets and are therefore not optimal; the lower threshold of approximately $p_{X}=10\%$ and consistently higher logical error rates (at $p_{X}=10\%$, the MWPM exhibits a logical error rate of approximately $26\%$ compared to the SM Monte Carlo $24-25\%$, for example) reflect this. This discrepancy in MWPM threshold from the $10.5(2)\%$ zero-temperature critical value found in Ref. [67] may also be accounted for by finite size effects.
@@ IV.2 Circuit compilations
The stabilizers of a QEC code can compiled into syndrome-extraction circuits in a variety of ways. The choice of implementation may be influenced by factors such as the native hardware-efficient gate set or a need to tolerate fabrication defects [35, 123, 84, 43]. In this section, we demonstrate how different syndrome-extraction circuits for the repetition code can be mapped onto spin diagrams with distinct bond anisotropies. We compare these models to offer insight into the differing performance of these circuits.
@@
McEwen et al. [84] introduced an alternate, dynamic syndrome-extraction circuit for the repetition code that exchanges the roles of data and ancilla qubits, producing a “wiggling” pattern in the detector cells. Figure 8b shows this circuit and the corresponding spin diagrams. As with the regular circuit, the simplified model (cf. Fig. 8d) is a random bond Ising model, this time on an un-sheared honeycomb lattice with anisotropic weights $w_{\braket{i,j}}\in\{2,2,3\}$. In the bulk of the model, the shear does not affect local interactions—the difference in memory-experiment threshold between these models should be solely due to the differing anisotropy.
Figure 9: MWPM error rates for memory experiments of the standard (solid line) and wiggling (dot-dashed line) repetition code circuits with N=2d1N=2d-1, T=8d1T=8d-1. The wiggling circuit has a lower threshold at $2.87(1)\%$ compared to $2.94(1)\%$ for the standard circuit.
To understand the difference in decoding performance of these two circuits, consider the free energy cost of inserting a spatial logical operator. One representative of this logical operator is the product of $X$ over all qubits in-between each measurement and reset, indicated by the circular errors and dashed bonds in Fig. 8. Starting from a ground state with $\eta_{X,\braket{i,j}}=1$ and $\sigma_{i}=1$ for both models, applying the logical error flips vertical weight-$3$ bonds. The energy cost is identical for both circuits. However, we need to also account for the contributions to the free energy from gauge-equivalent errors. Flipping a spin attached to the affected bonds (i.e., including a gauge operator to deform the logical error) changes energy by $\Delta\mathcal{E}_{\text{standard}}=2K_{X}^{(1)}$ in the standard circuit, and $\Delta\mathcal{E}_{\text{wiggling}}=4K_{X}^{(2)}-2K_{X}^{(3)}$ in the wiggling circuit. Taking the low-error limit of Eq. (25), one can show that
@@
We performed simulations of MWPM decoding for a memory experiment with both circuits. We simulated a circuit depth of T=8d1T=8d-1; this equates to $2d$ rounds of syndrome measurements, and initialization and final readouts. We plot the results of the memory experiment in Fig. 9: the standard circuit consistently achieves lower logical error rates, as we expected based on our spin models. The standard circuit has a MWPM threshold of approximately $2.94(1)\%$, while the wiggling circuit has a lower threshold of $2.87(1)\%$. This is also consistent with our expectations. Running Monte Carlo simulations of the spin models and estimating the ML decoding error rates, Fig. 10 shows a threshold of $3.03(1)\%$ for the standard circuit and $2.96(1)\%$ for the wiggling circuit, both slightly exceeding their MWPM thresholds. The logical error rates themselves are also lower for the ML decoder, achieving a rate of $21.5\%$ at $p_{X}=2.7\%$ with the $d=9$ standard circuit, compared to $23\%$ for the MWPM decoder, for example.
@@
Figure 10: Monte Carlo simulations of the ML logical error rate for two repetition code circuits with N=2d1N=2d-1, T=8d1T=8d-1. The standard circuit again exhibits a higher threshold and lower logical error rate compared to the wiggling circuit. For the standard circuit, a threshold of $3.03(1)\%$ exceeds the MWPM threshold of $2.94(1)\%$ from Fig. 9.